If "Y" has a geometric distribution with success probability 0.3,what is the largest value of "y",such that "P9Y>y)>0.1" ?


1
Expert's answer
2021-05-20T08:30:09-0400

Given "p=0.3"

Then "q=1-p=1-0.3=0.7"

By using the result "P(y>y_0)=q^{y_0}", we have


"P(y>y_0)=0.7^{y_0}"

If "P(y>y_0)>0.1," then


"0.7^{y_0}>0.1"

Let "0.7^{y_0}=0.1"

"\ln{0.7^{y_0}}=\ln{0.1}"

"y_0\ln{0.7}=\ln{0.1}"

"y_0=\dfrac{\ln 0.1}{\ln 0.7}\approx6.4557"

"0.7^{6}=0.117649>0.1"

"0.7^{7}=0.0823543<0.1"

The largest value of "y" is 6.